Score: 0%, 0 of 10 pts 8 of 9 (0 complete) HY Score: 0 of 2 pts P4-41A (similar to EQuestion Help Southwick Products manufactures its products in two separate departments: Machining and Assembly. Total manufacturing averhead costs for the year are budgeted at s 1,100,000, or his nout te Mach nig Departmert r an 9680,00° pmna iy for machine operaton and dopreciation) whi·tho Assen bly Department nars 5420,000. The compary estimates that it will inour 10,000 machine hours (all in the Machining Department) and 22,000 direct labor hours (8,000 in the Machining Department and 14,000 in the Assembly Department) during the year Southwick Products currenitly uses a plantwide qverhead rate based on direct labor hours to allocate overhead. However, the company is considering refining its overhead allocation system by using departmental overhead rates. The Machining Department would allocabe its overhead using machine hours (MH but the Assembly Department would allocate its overhead using direct labor (DL) hours. The following chart shows the machine hours (MH) and direct labor (DL) hours ncurred by Jobs 500 and 501 in each production department Read the rssuirements Requirement 1.Compute the company's oument plantwide overhead rate.(Round your answer to the nearest dollar) Begin by Cetermining the formula, then compute the rate Plantwide overhead rate Choose from any drop-down ist and then click Check Answer 12%.arrg Clear A Check Answer To see what to study next go to your Study Plan 5 8
s its products in two separate departments: Machining and Assembly. Total manufacturing overhead costs Machining Department inours $680,000 (primarily for machine operation and depreciation) while the Assemb es that it will incur 10,000 machine hours (all in the Machining Department) and 22,000 direct labor hours (8, ssembly Depactment)i during the vear es a plantwideData Table sing departme cate its overh the company is cons d using machine hou ours (MH) and direct each product dditional infom Machining Assembly Department Department 12 DL hours Job 500... .6 ompany's curn la, then compu MH 5 DL hours MH 5 DL hours Job 501... .12 12 DL hours Both Jobs 500 and 501 used $1,800 of direct materials. Wages and benefits total $30 per direct labor hour. Southwick Products prices its products at 110% of total manufacturing costs. Print Done n list and then click Check Answer. Clear All Check A tudy next, go to your Study Plan. 20
anufactures its products in two soparate departments: Machining and Assembly. Total manufacturing overhead costs fort ount, the Machining Department incurs $680,000 (primarily for machine operation and depreciation) while the ny Requirements and 1. Compute the company's current plantwide overhead rate viev 2. Compute refined departmental overhead rates. s. 3. Which job (Job 500 or Job 501) uses more of the company's resources? Explain. 4. Compute the total amount of overhead allocated to each job if the company uses its current plantwide overhead rate. 5. Compute the total amount of overhead allocated to each job if the company uses departmental theoverhead rates. 6. Do both allocation systems accurately reflect the resources that each job used? Explain. 7. Compute the total manufacturing cost and sales price of each job using the company's current plantwide overhead rate 8. Based on the current (plantwide) allocation system, how much profit did the company think it earned on each job? Based on the departmental overhead rates and the sales price determined in Requirement 7, how much profit did it really earn on each job? 9. Compare and comment on the results you obtained in Requirements 7 and &. Print Done Clear All Check Ans hat to study next, go to your Study Plan. F4 FS F7 F8 F9
Question Help E4-29A (similar to) Gooding idstres daduros freezers. Here is some infrmation ubod ED(Cilick the loon to view the information.) Read the teauiremens 17. line of freezers over to past year. otal COQ. (Round the percentage to the one tenth of a percent, XXS. If a box is not used in the table leave the box emply, do not select a label or enter a zero,) Total Costs of Percentage of Total Costs of Quality (Rounded) Gooding Cost of Quality Report Costs Prevention Costs Total prevention costs Appraisal Costs: otal appraisal cost Choose trom any list or enter any number in the input felds and then click Check Answer Clear All Check Answer To see what to study next go to your Study Plan 3 8
o view Data Table epare a the pen Number of units manufactured and sold abel or e 1,250 450 1,000 1.0 hours Cost of Variable manufacturing and seling costs per unit Hours spent by engineers improving design Testing and inspection hours per unit.. Percentage of units reworked in the plant after inspection. Average variable rework costs per refrigerator Percentage of units repaired under warranty at customers' homes verage variable repair costs at customers' homes per refrigerator. Estimated lost sales.due to quality reputation Design engineer's labor rate (includes benefits and payroill taxes).... Inspection costs per hour $ 520 10% $ 540 250 units $120 per hour $20 per hour costs .... osts Print Done ist or er Clear All Check Answer what to study next, go to your Study Plan F7 F8
ew the information.) centage re a Cost of Requirements t of Quality 1. Prepare a Cost of Quality (COQ) report. In addition to calculating the total costs 2. Why would the variable repair cost at the customers home generally be higher 3. of each category, calculate the percentage of COQ category of the total Coa. than the "in factory" rework cost? How can use of the COQ report lead to improvements in the production process or in the quality of the product itself? Print Done or enter any number in the input fields and then click Check Answer Clear All Check A t to study next, go to your Study Plan F4 F5 F6 F7 F9
Answer #1
1 PLANT WIDE OVERHEAD RATE
Plant wide Overhead Rate=Total Budgeted Overhead/Total Direct Labor hour
A Total budgeted manufacturing overhead cost $1,100,000
B Total budgeted direct labor hours          22,000
C=A/B Plant wide Overhead Rate $50 per labor hour
2 REFINED DEPARTMENTAL OVERHEAD RATES
Machining Department:
D Machining Department Overhead $680,000
E Budgeted machine hours          10,000
G=D/E Machining Department Overhead Rate $68 per machine hour
Assembly Department:
H Assembly Department Overhead $420,000
I Budgeted labor hours in assembly department          14,000
J=H/I Assembly Department Overhead Rate $30.00 per labor hour
3 Job 500 job 501
K Machine hours used 6 12
L Total Machine hour available          10,000                       10,000
M=L/K Total Capacity in number of pieces             1,667                             833
N Direct labor hour used 17 17
P Direct labor hour available          22,000                       22,000
Q=P/N Total Capacity in number of pieces       1,294.12                    1,294.12
Availability of machine hour is the constraint for increasing production.
This resource is limiting the capacity
Job 501 uses more of the critical resources of machining
Direct labor hour is not a critical resource ,since it can be increased without capital expenditure
4 Overhead allocated using Plant wide overhead rate
Overhead allocated to Job 500 $850 (17*$50)
Overhead allocated to Job 501 $850 (17*$50)
5 Overhead allocated using Departmental overhead rate
Overhead allocated to Job 500 $918 (6*68+17*30)
Overhead allocated to Job 501 $1,326 (12*68+17*30)
6 NO, Departmental allocation reflect the resources the job uses accurately
7 Based on Current Plant wide overhead Rate
Job500 Job501
A Direct material Costs $1,800 $1,800
B Direct Labor hours 17 17
C=B*30 Wages and benefits $510 $510
D=B*50 Overhead $850 $850
E=A+C+D Total Manufacturing costs $3,160 $3,160
F=110%*E Sales Price $3,476 $3,476
G=F-E Profit $316 $316
8 Based on Departmental Overhead rate
Job500 Job501
A Direct material Costs $1,800 $1,800
B Direct Labor hours 17 17
C=B*30 Wages and benefits $510 $510
D Overhead $918 $1,326
E=A+C+D Total Manufacturing costs $3,228 $3,636
F=110%*E Sales Price $3,551 $4,000
G=F-E Profit $323 $364